# On Function-Correcting Codes in the Lee Metric

### v2

@@ Abstract
Function-correcting codes are a coding framework designed to minimize redundancy while ensuring that specific functions or computations of encoded data can be reliably recovered, even in the presence of errors. The choice of metric is crucial in designing such codes, as it determines which computations must be protected and how errors are measured and corrected. Previous work by Liu and Liu [6] studied function-correcting codes over $\mathbb{Z}_{2^{l}},\ l\geq 2$ using the homogeneous metric, which coincides with the Lee metric over $\mathbb{Z}_{4}$. In this paper, we extend the study to codes over $\mathbb{Z}_{m},$ for any positive integer $m\geq 2$ under the Lee metric and aim to determine their optimal redundancy. To achieve this, we introduce irregular Lee distance codes and derive upper and lower bounds on the optimal redundancy by characterizing the shortest possible length of such codes. These general bounds are then simplified and applied to specific classes of functions, including Lee-local functions, Lee weight functions, and Lee weight distribution functions. We extend the bounds established by Liu and Liu [6] for codes over $\mathbb{Z}_{4}$ in the Lee metric to the more general setting of $\mathbb{Z}_{m}$. Additionally, we explicitly derive a Plotkin-like bound for linear function-correcting codes in the Lee metric. As the Lee metric coincides with the Hamming metric over the binary field, we demonstrate that our bound naturally reduces to a Plotkin-type bound for function-correcting codes under the Hamming metric over $\mathbb{Z}_{2}$. Furthermore, when the underlying function is bijective, function-correcting codes reduce to classical error-correcting codes. In parallel, our bound correspondingly reduces to the classical Plotkin bound for error-correcting codes, both for the Lee metric over $\mathbb{Z}_{m}$ and for the Hamming metric over $\mathbb{Z}_{2}$.
@@ 2 Preliminaries
[4, Lee distance metric] For $u=(u_{1},u_{2},\dots,u_{k}),v=(v_{1},v_{2},\dots,v_{k})\in\mathbb{Z}_{m}^{k}$, the Lee distance between $u$ and $v$ is defined as
@@
The set $\mathbb{Z}_{m}^{k}$ is called the message set. An encoding function $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+r}$ defined as $\mathcal{E}(u)=(u,p(u))$, where $p(u)\in\mathbb{Z}_{m}^{r}$ is the redundancy vector. The set
@@
[10, Proposition 10.10] Let $\rho<m/2$. Then the size of a ball in $\mathbb{Z}_{m}^{r}$ is given by
@@ 3 Function-correcting Lee metric codes
In error-correcting codes, we add redundancy to each message to enable the receiver to detect and correct any errors that affect the entire message during transmission. In contrast, in function-correcting codes, we need to correct messages with different function values. It allows for more efficient encoding. This approach reduces redundancy when several messages share the same function value. As a result, function-correcting codes offer a targeted and resource-efficient alternative to classical error correction. In this section, we formally define the function-correcting Lee metric code (in short, FCLMC) and establish an equivalence between the function-correcting Lee metric code and irregular-distance codes in the Lee metric. We discuss relations and bounds on the optimal redundancy of an $(f,t)$-FCLMC. Throughout the paper, we assume that a codeword $\mathcal{E}(u)$ is transmitted through a noisy channel, the received word is $a\in\mathbb{Z}_{m}^{k+r}$ with $d_{L}(\mathcal{E}(u),a)\leq t$, that is, at most $t$ errors occur during transmission. We define function-correcting Lee metric codes as follows.
@@
An encoding function $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+r}$ with $\mathcal{E}(u)=(u,p(u))$ defines a function-correcting Lee metric code with redundancy $r$ if $d_{L}(\mathcal{E}(u),\mathcal{E}(v))\geq 2t+1$ holds for all $u,v\in\mathbb{Z}_{m}^{k}$ with $f(u)\neq f(v)$.
@@
Let $u_{1},u_{2},\dots,u_{M}\in\mathbb{Z}_{m}^{k}$. Define the distance requirement matrix $D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M})$ of order $N\times N$ in the Lee metric for a given function $f$ as follows:
@@
Let $D$ be an $M\times M$ square matrix over positive integers. A code $C=\{C_{1},C_{2},\dots,C_{M}\}\subseteq\mathbb{Z}_{m}^{r}$ is a called $D$-code in Lee metric if $d_{L}(C_{i},C_{j})\geq[D]_{ij}$ for all $1\leq i,j\leq M$.
@@
Let $\{u_{1},u_{2},\dots u_{M}\}\subseteq\mathbb{Z}_{m}^{k}$ with $M\leq m^{k}$. Then the optimal redundancy of an $(f,t)$-FCLMC satisfies
@@
Let $N(D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta}))=r$, $D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})=D$ and $C=\{C_{i}\mid 1\leq i\leq\eta\}\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code of length $r$. Define an encoding function $\mathcal{E}(u)=(u,C_{f(u)})$. We claim this encoding defines an $(f,t)$-FCLMC. For this, let $u,v\in\mathbb{Z}_{m}^{k}$ with $e_{i}=f(u)\neq f(v)=e_{j}$. This implies that $C_{f(u)}\neq C_{f(v)}$ and
@@ 4 Bounds on the shortest length of irregular-distance codes
In this section, we present some bounds on $N_{L}(D)$ that give bounds on the optimal redundancy using results from the previous section. In the next lemma, we give a lower bound on $N_{L}(D)$ for a given Lee distance matrix $D$ which is a generalization of the Plotkin bound for codes with irregular Lee distance requirements.
@@
Let $D$ be an $M\times M$ Lee distance matrix. Then
@@
Let $N_{L}(D)=r$. Let $C\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code of length $r$ and size $M$. Consider an $M\times r$ matrix $A$ whose rows are codewords $\{C_{1},C_{2},\dots,C_{M}\}=C$. By [15, Eq. (15)], we have
@@
[6] Let $m=4$. For any Lee distance matrix $D\in\mathbb{N}_{0}^{M\times M}$, we have
@@
[5] Let $m=2$. For any Lee distance matrix $D\in\mathbb{N}_{0}^{M\times M}$, we have
@@
We construct a $D$-code of length $r$. First, select an arbitrary codeword $C_{1}$. For $j=2$, since $m^{r}>V_{L}(r,[D]_{12}-1)$, there exists $C_{2}\in\mathbb{Z}_{m}^{r}$ such that $d_{L}(C_{1},C_{2})\geq[D]_{12}$. For $j=3$, $m^{r}>V_{L}(r,[D]_{13}-1)+V_{L}(r,[D]_{23}-1)$, there exists $C_{3}\in\mathbb{Z}_{m}^{r}$ such that $d_{L}(C_{1},C_{3})\geq[D]_{13}$ and $d_{L}(C_{2},C_{3})\geq[D]_{23}$. Similarly, we chose $C_{4},C_{5},\dots,C_{M}$ such that $d_{L}(C_{i},C_{j})\geq[D]_{ij}$ for $i<j$. Thus, if
@@
then there exist a $D$-code, namely the code $\{C_{1},C_{2},\dots,C_{M}\}$. This completes the proof. ∎
A minor computational error appears in [5, Lemma 4]. We state the corrected version here (see also [16]).
Lemma 4.5.
For any $M,d\in\mathbb{N}$ with $d\geq 10$ and $M\leq d^{2}$, we have
$N_{H}(M,d)\leq\frac{2d-2}{1-2\sqrt{ln(d)/d}},$
where $N_{H}(M,d)$ is the smallest length code such that the code is a $D$-code with respect to the Hamming distance.
Lemma 4.6.
Let $D$ be a distance matrix with $D]_{ij}=d$ for all $i\neq j$. Then $N_{L}(M,d)\leq N_{H}(M,d)$.
@@
The proof follows from the fact that $d_{L}(x,y)\geq d_{H}(x,y)$ for all $x,y\in\mathbb{Z}_{m}$, where $d_{H}(x,y)$ is the Hamming distance between $x$ and $y$. ∎
Using Lemma 4.5 and Lemma 4.6, we have the following.
Corollary 4.7.
For any $M,d\in\mathbb{N}$ with $d\geq 10$ and $M\leq d^{2}$, we have
$N_{L}(M,d)\leq\frac{2d-2}{1-2\sqrt{ln(d)/d}}.$
Let $N_{L}(\eta,d)=\min\{n\mid\exists\text{a code }C\subseteq\mathbb{Z}_{m}^{n}\text{ with }d_{L}(C)\geq 2t,|C|=\eta\}$. That is, $N_{L}(\eta,d)$ is the minimum possible length of a code over $\mathbb{Z}_{m}$ with size $\eta$ and the minimum Lee distance $d$. The following theorem gives lower and upper bounds on the optimal redundancy of FCLMCs.

### v3

@@ Abstract
Function-correcting codes are a coding framework designed to minimize redundancy while ensuring that specific functions or computations of encoded data can be reliably recovered, even in the presence of errors. The choice of metric is crucial in designing such codes, as it determines which computations must be protected and how errors are measured and corrected. Previous work by Liu and Liu [7] studied function-correcting codes over $\mathbb{Z}_{2^{l}},\ l\geq 2$ using the homogeneous metric, which coincides with the Lee metric over $\mathbb{Z}_{4}$. In this paper, we extend the study to codes over $\mathbb{Z}_{m},$ for any positive integer $m\geq 2$ under the Lee metric and aim to determine their optimal redundancy. To achieve this, we introduce irregular Lee distance codes and derive upper and lower bounds on the optimal redundancy by characterizing the shortest possible length of such codes. These general bounds are then simplified and applied to specific classes of functions, including locally bounded functions, Lee weight functions, and Lee weight distribution functions. We extend the bounds established by Liu and Liu [7] for codes over $\mathbb{Z}_{4}$ in the Lee metric to the more general setting of $\mathbb{Z}_{m}$. Moreover, we give explicit constructions of function-correcting codes in Lee metric. Additionally, we explicitly derive a Plotkin-like bound for linear function-correcting codes in the Lee metric. As the Lee metric coincides with the Hamming metric over the binary field, we demonstrate that our bound naturally reduces to a Plotkin-type bound for function-correcting codes under the Hamming metric over $\mathbb{Z}_{2}$. Furthermore, when the underlying function is bijective, function-correcting codes reduce to classical error-correcting codes. In parallel, our bound correspondingly reduces to the classical Plotkin bound for error-correcting codes, both for the Lee metric over $\mathbb{Z}_{m}$ and for the Hamming metric over $\mathbb{Z}_{2}$.
@@ 2 Preliminaries
[5, Lee distance metric] For $u=(u_{1},u_{2},\dots,u_{k}),v=(v_{1},v_{2},\dots,v_{k})\in\mathbb{Z}_{m}^{k}$, the Lee distance between $u$ and $v$ is defined as
@@
The set $\mathbb{Z}_{m}^{k}$ is called the message set. An encoding function $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+r}$ is defined as $\mathcal{E}(u)=(u,p(u))$, where $p(u)\in\mathbb{Z}_{m}^{r}$ is the redundancy vector. The set
@@
[13, Proposition 10.10] Let $\rho<m/2$. Then the size of a ball in $\mathbb{Z}_{m}^{r}$ is given by
@@
Let $r$ be a positive integer and $\rho$ be a real number. Denote
$B_{r}(\rho)=\{x\in\mathbb{Z}_{m}^{r}|\ w_{L}(x)\leq r\rho\}.$
In [8], Loeliger established an upper bound on $B_{r}(\rho)$ for any coordinate-additive metric, via the entropy of an auxiliary probability distribution. We state the bound for the Lee metric using [8, Eq.(5)].
Lemma 2.3.
Let $0<\rho\leq 1$ be a real number. Then
$|B_{r}(\rho)|\leq(e^{\lambda\rho}\mathcal{L}(\lambda))^{r},$
holds for all non-negative integer $\lambda$ and $r$, where $\mathcal{L}(\lambda)=\sum_{a\in\mathbb{Z}_{m}}e^{-w_{L}(a)}$ and $e$ is natural constant.
Lemma 2.4.
For $m\geq 5$, we have $\sum_{a\in\mathbb{Z}_{m}}e^{-w_{L}(a)}<3$. Moreover, for $m\leq 4$, $\sum_{a\in\mathbb{Z}_{m}}e^{-w_{L}(a)}\leq 2$.
Proof.
If $m=2\ell$ for some positive integer $\ell$, then
$\sum_{a\in\mathbb{Z}_{m}}e^{-w_{L}(a)}=e^{0}+e^{-\ell}+2\sum_{i=1}^{\ell-1}e^{-i}=1+e^{-\ell}+2(e^{-1}+e^{-2}+\dots+e^{-(\ell-1)})$
$=1+e^{-\ell}+2\frac{e^{-1}(1-e^{-\ell+1})}{1-e^{-1}}=1+e^{-\ell}+2\frac{1-e^{-\ell+1}}{e-1}$
$\leq 1+e^{-\ell}+2(1-e^{-\ell+1})=e^{-\ell}(1-2e)+3<3.$
Similarly, if $m=2\ell+1$ for some positive integer $\ell$, then
$\sum_{a\in\mathbb{Z}_{m}}e^{-w_{L}(a)}=e^{0}+2e^{-\ell}+2\sum_{i=1}^{\ell-1}e^{-i}=1+2e^{-\ell}+2(e^{-1}+e^{-2}+\dots+e^{-(\ell-1)})$
$=1+2e^{-\ell}+2\frac{e^{-1}(1-e^{-\ell+1})}{1-e^{-1}}=1+2e^{-\ell}+2\frac{1-e^{-\ell+1}}{e-1}$
$\leq 1+2e^{-\ell}+2(1-e^{-\ell+1})=2e^{-\ell}(1-e)+3<3.$
For $m\leq 4$, the bound is straightforward. ∎
@@ 3 Function-correcting Lee metric codes
In error-correcting codes, we add redundancy to each message to enable the receiver to detect and correct any errors that affect the entire message during transmission. In contrast, in function-correcting codes, we need to correct messages with different function values. It allows for more efficient encoding. This approach reduces redundancy when several messages share the same function value. As a result, function-correcting codes offer a targeted and resource-efficient alternative to classical error correction. In this section, we formally define the function-correcting Lee metric code (in short, FCLMC) and establish an equivalence between the function-correcting Lee metric code and irregular-distance codes in the Lee metric. We discuss relations and bounds on the optimal redundancy of an $(f,t)$-FCLMC. Throughout the paper, we assume that a codeword $\mathcal{E}(u)$ is transmitted through a noisy channel, the received word is $y\in\mathbb{Z}_{m}^{k+r}$ with $d_{L}(\mathcal{E}(u),y)\leq t$. In this case, when the Lee distance between the transmitted code vector and received vector is at most $t$, we say that at most $t$ errors occur during transmission, although it does not mean that $t$ coordinates changed during transmission. We define function-correcting Lee metric codes as follows.
@@
An encoding function $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+r}$ with $\mathcal{E}(u)=(u,p(u))$ defines an $(f,t)$-function-correcting Lee metric code (FCLMC) with redundancy $r$ if $d_{L}(\mathcal{E}(u),\mathcal{E}(v))\geq 2t+1$ holds for all $u,v\in\mathbb{Z}_{m}^{k}$ with $f(u)\neq f(v)$.
Remark 3.2.
The encoding algorithm for function-correcting codes involves a systematic encoding, that is, the encoded message consists of the first $k$ coordinates corresponding to the message vector, followed by the concatenation of $r$ coordinates of the redundancy vector. The process is done by an encoding function $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+r}$ with $\mathcal{E}(u)=(u,p(u))\in\mathbb{Z}_{m}^{k+r}$, where $p(u)$ is called the redundancy vector of length $r$. The collection $\{\mathcal{E}(u)|\ u\in\mathbb{Z}_{m}^{k}\}\subseteq\mathbb{Z}_{m}^{k+r}$ is called a code over $\mathbb{Z}_{m}$ of length $k+r$ (or equivalently, with redundancy $r$), and individual encoded messages $\mathcal{E}(u)$ are called codewords. Let $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+r}$ defines an $(f,t)$-FCLMC. For $u\in\mathbb{Z}_{m}^{k}$, $\mathcal{E}(u)\in\mathbb{Z}_{m}^{k+r}$ is transmitted through a noisy channel and receiver received $y\in\mathbb{Z}_{m}^{k+r}$. By Definition 3.1, the encoding $\mathcal{E}$ allows the receiver to determine the correct value of $f(u)$ whenever $d_{L}(\mathcal{E}(u),y)\leq t$ as follows.
Consider set $S_{y}=\{v\in\mathbb{Z}_{m}^{k}|\ d_{L}(\mathcal{E}(v),y)\leq t\}$. The set is non-empty as $u\in S_{y}$. Then by triangle inequality, we have
$d_{L}(\mathcal{E}(u),\mathcal{E}(v))\leq d_{L}(\mathcal{E}(u),y)+d_{L}(y,\mathcal{E}(v))\leq 2t,\forall v\in S_{y}.$
Thus, by Definition 3.1, $f(u)=f(v)$ for all $v\in S_{y}$. Therefore, to obtain correct value of $f(u)$, the receiver searches for a vector $v\in\mathbb{Z}_{m}^{k}$ with $d_{L}(\mathcal{E}(v),y)\leq t$. Once such a vector $v$ found, the receiver terminates the search and decodes $f(u)$ as $f(v)$. Note that although $u\in S_{y}$, the found vector $v$ need not coincide with $u$, that is, under these conditions, the receiver may not recover the exact message $u$.
@@
Let $u_{1},u_{2},\dots,u_{M}\in\mathbb{Z}_{m}^{k}$. Define the Lee distance requirement matrix $D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M})$ of order $M\times M$ in the Lee metric for a given function $f$ as follows:
@@
Let $D$ be an $M\times M$ square matrix over non-negative integers with diagonal entries zero. A code $C=\{C_{1},C_{2},\dots,C_{M}\}\subseteq\mathbb{Z}_{m}^{r}$ is a called $D$-code in Lee metric if $d_{L}(C_{i},C_{j})\geq[D]_{ij}$ for all $1\leq i,j\leq M$.
@@
Let $\{u_{1},u_{2},\dots u_{M}\}\subseteq\mathbb{Z}_{m}^{k}$ with $M\leq m^{k}$. Then the optimal redundancy of $(f,t)$-FCLMCs satisfies
@@
Let $N_{L}(D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta}))=r$, $D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})=D$ and $C=\{C_{e_{i}}\mid 1\leq i\leq\eta\}\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code in the Lee metric of length $r$. Define an encoding function $\mathcal{E}(u)=(u,C_{f(u)})$. We claim this encoding defines an $(f,t)$-FCLMC. For this, let $u,v\in\mathbb{Z}_{m}^{k}$ with $e_{i}=f(u)\neq f(v)=e_{j}$. This implies that $C_{f(u)}\neq C_{f(v)}$ and
@@ 4 Bounds on the shortest length of irregular-distance codes
In this section, we present some bounds on $N_{L}(D)$ that give bounds on the optimal redundancy using results from the previous section. In the next lemma, we give a lower bound on $N_{L}(D)$ for a matrix $D$ over non-negative integers, which is a generalization of the Plotkin bound for codes with irregular Lee distance requirements.
@@
Let $D$ be an $M\times M$ matrix over non-negative integers with diagonal entries zero. Then
@@
Let $N_{L}(D)=r$. Let $C\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code in the Lee metric of length $r$ and size $M$. Consider an $M\times r$ matrix $A$ whose rows are codewords $\{C_{1},C_{2},\dots,C_{M}\}=C$. By [18, Eq. (15)], we have
@@
[7] Let $m=4$. For any matrix $D\in\mathbb{N}_{0}^{M\times M}$ with diagonal entries zero, we have
@@
[6] Let $m=2$. For any matrix $D\in\mathbb{N}_{0}^{M\times M}$ with diagonal entries zero, we have
@@
We construct a $D$-code in the Lee metric of length $r$, where
$m^{r}\geq\max_{j\in[M]}\sum_{i=1}^{j-1}V_{L}(r,[D]_{ij}-1).$
First, select an arbitrary codeword $C_{1}$. For $j=2$, since $m^{r}>V_{L}(r,[D]_{12}-1)$, there exists $C_{2}\in\mathbb{Z}_{m}^{r}$ such that $d_{L}(C_{1},C_{2})\geq[D]_{12}$. For $j=3$, $m^{r}>V_{L}(r,[D]_{13}-1)+V_{L}(r,[D]_{23}-1)$, there exists $C_{3}\in\mathbb{Z}_{m}^{r}$ such that $d_{L}(C_{1},C_{3})\geq[D]_{13}$ and $d_{L}(C_{2},C_{3})\geq[D]_{23}$. Similarly, we chose $C_{4},C_{5},\dots,C_{M}$ such that $d_{L}(C_{i},C_{j})\geq[D]_{ij}$ for $i<j$. Thus, if
@@
then there exist a $D$-code in the Lee metric, namely the code $\{C_{1},C_{2},\dots,C_{M}\}$. This completes the proof. ∎
Recall that $N_{L}(M,d)$ is smallest possible length of a $D$-code in Lee metric, where $D\in\mathbb{N}_{0}^{M\times M}$ and $D]_{ij}=d$ for all $i\neq j$. Using Lemma 4.4 and Lemma 2.3, we have an upper bound on $N_{L}(M,d)$.
Corollary 4.5.
For any $M,d,m\in\mathbb{N}$ with $M>1,m\geq 5$, we have
$N_{L}(M,d)\leq\left\lceil\frac{\ln(M-1)+(d-1)}{\ln m-\ln 3}\right\rceil$
@@
Let $D$ be a matrix with $D]_{ij}=d$ for all $i\neq j$. Then by Lemma 4.4,
$N_{L}(M,d)\leq\min_{r\in\mathbb{N}}\left\{r:\ m^{r}>\max_{j\in[M]}\sum_{i=1}^{j-1}V_{L}(r,d-1)\right\}$
$\leq\min_{r\in\mathbb{N}}\left\{r:\ m^{r}>(M-1)V_{L}(r,d-1)\right\}.$
Take $\lambda=1$ in Lemma 2.3, we have
$V_{L}(r,d-1)=\left|B_{r}\left(\frac{d-1}{r}\right)\right|\leq e^{d-1}\left(\sum_{a\in\mathbb{Z}_{m}}e^{-w_{L}(a)}\right)^{r}<e^{d-1}3^{r}.$
Consequently, $m^{r}>(M-1)V_{L}(r,d-1)$ whenever $r>\frac{\ln(M-1)+(d-1)}{\ln m-\ln 3}$. Hence, $N_{L}(M,d)\leq\left\lceil\frac{\ln(M-1)+(d-1)}{\ln m-\ln 3}\right\rceil$. ∎
Remark 4.6.
The Lee metric coincides with the homogeneous metric for $m=4$ and with the Hamming metric for $m=2$. In [7] and [6], the authors derived bounds on $N_{L}(M,d)$ for $m=4$ and $m=2$, respectively.
Let $N_{L}(\eta,d)=\min\{n\mid\exists\text{a code }C\subseteq\mathbb{Z}_{m}^{n}\text{ with }d_{L}(C)\geq d,|C|=\eta\}$. That is, $N_{L}(\eta,d)$ is the minimum possible length of a code over $\mathbb{Z}_{m}$ with size $\eta$ and the minimum Lee distance $d$. The following theorem gives lower and upper bounds on the optimal redundancy of FCLMCs.
@@
Remark 4.8.
For any $f:\mathbb{Z}_{m}^{k}\to Im(f)$ with $Im(f)=\{e_{1},e_{2},\dots,e_{\eta}\}$, by Theorem 3.10, the optimal redundancy of FCLMCs is upper bounded by $N_{L}(D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta}))$. Note that
$[D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})]_{ij}=[2t+1-d_{L}(e_{i},e_{j})]^{+}\leq 2t\ \ \text{ for all }i\neq j.$
Let $C$ be the code constructed in the proof of Theorem 4.7. Since
$d_{L}(C_{i},C_{j})\geq 2t\geq[D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})]_{ij},$
$C$ is a $D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})$-code in Lee metric. Thus,
$N_{L}(D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta}))\leq N_{L}(\eta,2t).$
Hence, Theorem 3.10 gives a sharper bound than Theorem 4.7. Nevertheless, Theorem 4.7 remains significant as it gives an explicit bound together with an explicit construction (see Proposition 4.10) that depends only on the cardinality of the image set and $t$. While Theorem 3.10 requires first computing the Lee distances between distinct function values in order to construct the matrix $D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})$ and subsequently determining the value $N_{L}(D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta}))$.
Theorem 4.7 suggests that determining the value $N_{L}(2,2t)$ and $N_{L}(\eta,2t)$ will give lower and upper bounds on the optimal redundancy of $(f,t)$-FCLMCs.
Proposition 4.9.
For positive integers $t$ and $m\geq 2$, we have
Proof.
Thus, $N_{L}(2,2t)\leq s$.
Therefore, $N_{L}(2,2t)\geq s$. This completes the proof. ∎
In the following result, we construct a code over $\mathbb{Z}_{m}$ consisting of $\lambda$ codewords with minimum Lee distance $2t$ . This construction yields an upper bound on $N_{L}(\lambda,2t)$ for arbitrary positive integers $\lambda$ and $t$.
Proposition 4.10.
Let $\lambda$ and $t$ be positive integers. Then
$N_{L}(\lambda,2t)\leq\lambda\left\lceil\frac{t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil.$
Proof.
Let $n=\left\lceil\frac{t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil$, and $\bm{0}=(0,0,\dots,0)$, $\bm{a}=\left(\left\lfloor\frac{m}{2}\right\rfloor,\left\lfloor\frac{m}{2}\right\rfloor,\dots,\left\lfloor\frac{m}{2}\right\rfloor\right)\in\mathbb{Z}_{m}^{n}$. Consider code $C=\{C_{1},C_{2},\dots,C_{\lambda}\}\subseteq\mathbb{Z}_{m}^{\lambda n}$, where for $1\leq i\leq\lambda$ is given by
$C_{i}=(\bm{0},\dots,\bm{0},\underbrace{\bm{a}}_{i\text{-th-place}},\bm{0},\dots,\bm{0})\in(\mathbb{Z}_{m}^{n})^{\lambda}.$
Observe that for $i\neq j$,
$d_{L}(C_{i},C_{j})=2w_{L}(\bm{a})$ $=2n\left\lfloor\frac{m}{2}\right\rfloor=2\left\lceil\frac{t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil\left\lfloor\frac{m}{2}\right\rfloor$
$\geq 2\frac{t}{\left\lfloor\frac{m}{2}\right\rfloor}\left\lfloor\frac{m}{2}\right\rfloor=2t,$
since $\lceil x\rceil\geq x$. Thus $C$ has the minimum Lee distance $2t$. This completes the proof. ∎
Remark 4.11.